OpenStudy (anonymous):

The manager of a large apartment complex knows from experience that 100 units will be occupied if the rent is 350 dollars per month. A market survey suggests that, on the average, one additional unit will remain vacant for each 5 dollar increase in rent. Similarly, one additional unit will be occupied for each 5 dollar decrease in rent. What rent should the manager charge to maximize revenue?

OpenStudy (cruffo):

let x = # of $5 price decreases. Revenue = Price times Quantity where Price = 350 - 5x and Quantity = 100 + 1x

OpenStudy (anonymous):

so what will equal?

OpenStudy (cruffo):

your back..

OpenStudy (anonymous):

yeah sorry had to run an errand

OpenStudy (cruffo):

np... do you understand why I'm letting x be the number of price decreases?

OpenStudy (anonymous):

becuase when the price is lower more people will rent?

OpenStudy (cruffo):

basically. It's one of those "tricks" you pick up :)

OpenStudy (cruffo):

So now onto the revenue function: \(\large R(x) = (350 - 5x)(100+x)\) FOIL to get \(\large R(x) = 35000-150x - 5x^2\)

OpenStudy (anonymous):

thats what i got, but im confused on what to do next

OpenStudy (cruffo):

What rent should the manager charge to maximize revenue? Key words are "maximize" and "revenue" We have made the revenue function. Now to "maximize". That means take the derivative, set it =0, solve for x.

OpenStudy (anonymous):

-150+10x=0?

OpenStudy (anonymous):

-150-10x=0?

OpenStudy (cruffo):

Good.

OpenStudy (anonymous):

and itll be 15

OpenStudy (cruffo):

actually -15

OpenStudy (anonymous):

ohh and then you subtract that from 350?

OpenStudy (cruffo):

Hmmmm. Since x was the number of price decreases ... so x = 1 would mean subrtact $5 from the original rent, x = 2 would mean subtract $10 from the original rent, But x = -15 means we should actually increase the rent by 15*5 = $75 So 350 + 75 = 425

OpenStudy (anonymous):

oohhh i understand

OpenStudy (anonymous):

thank you so much

OpenStudy (cruffo):

np

OpenStudy (anonymous):

do you know how to find the value of marginal profit?

OpenStudy (anonymous):

can you explain to me The demand function and the cost function for units of a product are and . Find value of the marginal profit when 8100. The value of the marginal profit when 8100 is